# Deep Reinforcement Learning for Robotic Bipedal Locomotion: A Brief Survey

### v5

@@ IV-B Reality gap: From simulation to the physical world
Another challenge hindering the deployment of DRL policies on bipedal robots is the simulation-to-reality (sim-to-real) gap. This refers to the significant discrepancy between a policy’s performance in a physics simulator and its performance on actual hardware. This gap is a critical obstacle because training directly on physical robots is often impractical. The millions of environmental interactions required for DRL would lead to accelerated mechanical wear, risk of catastrophic failure, and require constant human supervision. While simulation offers a safe and efficient alternative, the ultimate goal of “zero-shot” transfer, where a policy works perfectly without any real-world fine-tuning, is rarely achieved.
@@
A large body of research validates impressive locomotion skills purely within simulation, without attempting transfer to a physical system [25, 26, 72, 73, 34]. Even when transfer is successful, it often comes with compromises. Many successful transfers are not truly “zero-shot” and rely on a subsequent phase of extensive real-world fine-tuning or manual parameter tuning [74, 46]. In cases where policies do transfer without fine-tuning, they often exhibit a noticeable degradation in performance, where the robustness and agility seen in simulation are significantly lower in the real world [49, 33, 75].
@@
  • Robot dynamics modelling and actuation: Simulators struggle to perfectly replicate the complex dynamics of a physical bipedal robot, whose inherent instability makes it particularly sensitive to modeling errors. Factors such as motor friction, gear backlash, and precise link inertia are often simplified.
  • Contact and terrain modelling: Accurately simulating the intermittent foot-ground contact that bipeds rely on is extremely difficult. A mismatch between simulated and real-world friction or surface properties can cause unexpected slips or bounces, leading to a loss of balance.
  • Sensing and state estimation: A simulated robot has access to perfect, noise-free state information. In the real world, these states must be estimated from noisy sensors like IMUs and joint encoders [46, 74]. For a bipedal robot, precise state estimation is critical for maintaining balance.
Simulators like Isaac Gym [67], RoboCup3D [76], OpenAI Gym [77], and MuJoCo [78] detailed in Appendix B are widely used to train policies that closely mimic real-world physical conditions. These platforms use full-order dynamics to better represent the complex interactions robots face, and numerous sim-to-real frameworks [46, 79, 80] have demonstrated efficient and high-performance results. Despite these advancements, a significant gap persists between simulation and reality, exacerbated by the approximations made in simulation and the unpredictability of physical environments.
@@ IV-C Safety-critical Locomotion
Beyond performance metrics such as agility and robustness, the practical deployment of bipedal robots in human-centric environments is fundamentally contingent upon safety [1, 15, 16]. This includes ensuring the robot’s own integrity to prevent costly damage, as well as guaranteeing the safety of the surrounding environment and any humans within it. While many existing frameworks have demonstrated impressive locomotion skills, they often prioritize performance over these safety considerations. This creates a critical barrier that separates success in controlled laboratory settings from reliable operation in the unpredictable real world.
@@
  • Blind locomotion policies: Many current frameworks rely solely on internal sensors (proprioception) like joint angles and IMU data [47, 49, 32], creating a major safety risk. Lacking external perception, these robots cannot anticipate obstacles, slopes, or slippery surfaces, making them purely reactive and highly prone to failure. Despite these significant safety drawbacks, this approach is often adopted for several key reasons. Omitting vision simplifies the control problem to pure motor skills and avoids the computational complexity of real-time visual processing. Furthermore, since robust blind locomotion has already been demonstrated, vision is often treated as a supplementary component used to enhance task-specific precision [55] or path planning [60], rather than a core requirement for basic stability.
@@
  • Lack of physical constraint satisfaction: Many DRL frameworks lack built-in mechanisms to guarantee physical constraint satisfaction. This gap has motivated constrained or safety-aware DRL that enforces limits via the learning objective or auxiliary safety modules, for example Safe-RL on humanoids [81], hybrid DRL with identified low-dimensional safety models [82], footstep-constrained DRL policies [54], and reactive DRL steppers operating under feasibility constraints on uneven terrain [72]. This limitation makes it difficult to prevent the robot from exceeding joint limits, applying excessive torques, or causing self-collisions, particularly when reacting to unexpected events. This is a key area where constrained RL could be applied.
@@
In summary, the pursuit of performance in DRL has often sidelined critical safety issues. The prevalence of blind policies that cannot anticipate environmental hazards, combined with the lack of inherent mechanisms to enforce physical constraints, creates a significant risk that hinders real-world deployment. While these challenges are significant, they also define a clear path forward. The following section on Future Directions and Opportunities will explore specific research avenues, such as vision-based learning and safe reinforcement learning, aimed directly at overcoming these safety barriers and enabling the development of truly robust and reliable bipedal robots.
@@ V-A1 Multi-skill learning
A fundamental goal for the next generation of bipedal robots is to move beyond the paradigm of single-task specialisation and towards versatile skill learning [63, 83, 35]. This research direction focuses on enabling robots to acquire, adapt, and deploy a broad and varied repertoire of motor skills, allowing them to handle unforeseen situations and operate effectively in unstructured environments.
To achieve such versatility, researchers are pursuing several HL pathways, which can be broadly categorized into structured and holistic approaches. The structured approach focuses on explicit decomposition. A prominent example is hierarchical learning, where success depends on appropriately dividing responsibilities; for instance, a HL planner generates reference trajectories, while a LL DRL controller executes them robustly [59, 56], has shown in Section III. Similarly, skill composition employs a supervisor policy to select and sequence LL experts to solve complex tasks [26]. A related technique, knowledge distillation, leverages experts by first training them and then distilling their capabilities into a single compact generalist policy [84].
@@ V-A2 Traversing challenging environments
The goal of versatile skill learning is to enable bipedal robots to traverse challenging, human-centric environments where their unique form has an advantage. Validating capabilities on such terrains serves a crucial dual purpose. It tests a policy’s generalisation across diverse settings, including stairs and uneven ground, which is essential for real-world integration [49, 52, 63]. More critically, it benchmarks precision on treacherous paths such as stepping stones that demand exact foot placement [33, 19, 72]. These environments are the ultimate test of both a robot’s skill repertoire and its control fidelity.
@@ V-A3 Efficient learning and reward shaping
As detailed in Section IV, while DRL has unlocked impressive capabilities in bipedal locomotion, its reliance on training from scratch leads to significant sample inefficiency [40, 32]. Addressing this bottleneck is a crucial research frontier that calls for both more efficient algorithms and more robust reward designs.
To mitigate sample inefficiency for complex skills, several research pathways are being actively explored. A primary strategy is to leverage prior data rather than learning entirely from scratch. Leveraging prior knowledge provides strong guidance and reduce unsafe exploration by anchoring policies to feasible motion patterns [26, 45, 59]. Curriculum learning further organizes training from simple to progressively harder tasks, for example standing and balancing before walking and running, which improves stability and convergence [53, 51, 39].
Complementing advances in algorithms is the design of effective and robust rewards. Manual reward engineering remains a significant obstacle, since small choices can induce reward hacking and lengthy tuning cycles [68, 26]. Phase aware objectives are well established for cyclic gaits such as walking [32], whereas reward design for nonperiodic skills such as jumping is less standardized and often task specific [48]. Promising directions reduce manual effort by adding higher level guidance, including event based terms, goal conditioned objectives, and kinematic reference tracking [54, 28, 50]. Alternatively, learning rewards from data through inverse methods and related approaches aims to replace hand crafted objectives with implicit ones inferred from demonstrations [85]. Together, these directions seek to minimize skill specific tuning and improve the transferability and reliability of learned locomotion policies.
@@ V-A4 Motion retargeting
As human-like agents, bipedal robots, especially humanoids, have the unique advantage of a morphology that is similar to our own. This presents a significant opportunity: the potential to learn from vast libraries of human motion data. While large-scale datasets like AMASS [86] and Motion-X [87] provide a wealth of such data, they are inherently human-centric and cannot be used directly, requiring substantial retargeting effort [29]. Therefore, Motion Retargeting emerges as a critical component to bridge this gap. The challenge of this pathway is not merely to transfer human movements to the robot, but to generate trajectories that are both high-quality in their stylistic fidelity and rigorously physically feasible, adhering to the robot’s unique dynamics and constraints. Successfully developing these retargeting methods provides a scalable solution for accessing the data needed to train the natural and versatile generalist policies of the future.
@@ V-A5 Bridging the gap from simulation to reality
Strategies to bridge the sim-to-real gap generally follow two main philosophies. The first aims to train policies that are robust enough to tolerate the inevitable mismatch between simulation and reality. The second focuses on minimizing the gap itself by making the simulator a more faithful replica of the physical world.
The first approach seeks to reduce the discrepancy by improving the simulation’s fidelity. This is often achieved through system identification (SI), where real-world robot data is used to fine-tune the simulator’s parameters to create a more accurate ”digital twin” [74, 88]. This can include explicitly learning complex actuator dynamics to model the motors’ behaviour [74, 89]. Other methods, such as designing specialized feedback controllers [90], also contribute by making the system less sensitive to the remaining modeling errors.
In contrast, the second philosophy accepts that simulations will always be imperfect and instead focuses on creating highly adaptive, robust policies. The primary method here is DR, which forces a policy to generalize by training it across a wide range of simulated physical variations. Other various ways, such as through end-to-end training that uses measurement histories to adapt online like in RMA [79], or via policy distillation, where a privileged “teacher” guides a “student” policy [52] to have a knowledge of unknown information like friction . Additionally, techniques like adversarial motion priors [31, 30] are used to ensure the learned behaviours are not just robust but also physically plausible.
@@ V-A6 Perception-conditioned locomotion
ntegrating exteroceptive sensors like cameras and LiDAR enables bipedal robots to proactively plan footsteps, avoid obstacles, and adapt to upcoming terrain. This shift from reactive to anticipatory control is essential for navigating unstructured real-world environments.
The vision-based pathway is a human-inspired approach using RGB and depth cameras to capture rich data on colour, texture, and object appearance [33, 55, 39]. In contrast, LiDAR is an active sensing method that generates precise 3D point clouds of the terrain. While vision provides richer data but is sensitive to lighting, LiDAR offers robust geometric measurements without visual detail.
Based on this sensory data, current research is exploring two primary pathways for processing perceptual information for control. The first involves creating an intermediate geometric representation, such as a height map from scanners [55]. This provides the policy with structured topographical data for effective footstep planning. The second is a more end-to-end approach, which utilizes direct vision inputs like RGB or depth images as inputs to the RL policy for real-time decision-making [60, 91]. The former offers interpretability, while the latter promises more nuanced, reactive behaviours learned directly from raw perception.

### v6

@@ IV-B Challenges in transferring from simulation to reality
Another challenge hindering the deployment of DRL policies on bipedal robots is the sim-to-real gap. This refers to the significant discrepancy between a policy’s performance in a physics simulator and its performance on actual hardware. This gap is a critical obstacle because training directly on physical robots is often impractical. The millions of environmental interactions required for DRL would lead to accelerated mechanical wear, a risk of catastrophic failure, and require constant human supervision. While simulation offers a safe and efficient alternative, the ultimate goal of “zero-shot” transfer, where a policy works perfectly without any real-world fine-tuning, is rarely achieved.
@@
A large body of research validates impressive locomotion skills purely within simulation, without attempting transfer to a physical system [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2021_DeepQ_DNQ_stepping_animationlonly_steppingstones, 2017_Josh_drl_animation_imitationlearning_GRAIL_TRPO_MUJOCO, 2023_Chongben_jumping_DDPG_multi-task_network_simulationonly]. Even when transfer is successful, it often comes with compromises. Many successful transfers are not truly “zero-shot” and rely on a subsequent phase of extensive real-world fine-tuning or manual parameter tuning [2019_Yu_Phy_system_identification, 2020_Xie_firstsim2real_]. In cases where policies do transfer without fine-tuning, they often exhibit a noticeable degradation in performance, where the robustness and agility seen in simulation are significantly lower in the real world [2021_siekmann_blind_DRL_stair, 2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2020_park_recoverstability_drl_PPO_referencebased_simulation].
@@
  • Robot dynamics modelling and actuation: Simulators struggle to replicate the complex dynamics of a physical bipedal robot, whose inherent instability makes it particularly sensitive to modelling errors. Factors such as motor friction, gear backlash, and precise link inertia are often simplified.
  • Contact and terrain modelling: Accurately simulating intermittent foot–ground contact is extremely difficult. A mismatch between simulated and real-world friction or surface properties can cause unexpected slips or bounces, leading to loss of balance.
  • Sensing and state estimation: A simulated robot has access to perfect, noise-free state information. In the real world, these states must be estimated from noisy sensors such as IMUs and joint encoders [2020_Xie_firstsim2real_, 2019_Yu_Phy_system_identification]. For a bipedal robot, precise state estimation is critical for maintaining balance.
Simulators such as Isaac Gym [2021_Sim_Isaac], RoboCup3D [2003_Sim_Robocup], OpenAI Gym [2016_Sim_OpenAI], and MuJoCo [2012_Sim_Mujoco], detailed in Appendix B, are widely used to train policies that closely mimic real-world physical conditions. These platforms use full-order dynamics to better represent the complex interactions robots face, and numerous sim-to-real frameworks [2020_Xie_firstsim2real_, 2022_Ashish_RMA_sim2real_UCB, 2023_Rohan_sim2real_currentdecompose] have demonstrated efficient and high-performance results. Despite these advancements, a significant gap persists between simulation and reality, exacerbated by the approximations made in simulation and the unpredictability of physical environments.
@@ IV-C Safety-critical locomotion
Beyond performance metrics such as agility and robustness, the practical deployment of bipedal robots in human-centric environments is fundamentally contingent upon safety [2024_tong_advancements_review, 2021_Jenna_bipedrobot_review, 2021_Justin_review_recentporgressinleggedrobot]. This includes ensuring the robot’s own integrity to prevent costly damage, as well as guaranteeing the safety of the surrounding environment and any humans within it. While many existing frameworks have demonstrated impressive locomotion skills, they often prioritise performance over these safety considerations. This creates a critical barrier that separates success in controlled laboratory settings from reliable operation in the unpredictable real world.
@@
  • Blind locomotion policies: Many current frameworks rely solely on internal sensors (proprioception) such as joint angles and IMU data [2020_Siekmann_drl_e2e_s2r_cassie, 2021_siekmann_blind_DRL_stair, 2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie], creating a major safety risk. Lacking external perception, these robots cannot anticipate obstacles, slopes, or slippery surfaces, making them purely reactive and highly prone to failure. Despite these significant safety drawbacks, this approach is often adopted for several reasons: omitting vision simplifies the control problem to pure motor skills and avoids the computational cost of real-time visual processing. Moreover, since robust blind locomotion has already been demonstrated, vision is often treated as a supplementary component used to enhance task-specific precision [2023_duan_OSUDRL_heightmap_visionbased_hybrid] or path planning [2018_Tsunekawa_visualnavigation_hierarchy_DDPG-sim2real_NAO], rather than a core requirement for basic stability.
@@
  • Lack of physical constraint satisfaction: Many DRL frameworks lack built-in mechanisms to guarantee physical constraint satisfaction. This gap has motivated constrained or safety-aware DRL that enforces limits via the learning objective or auxiliary safety modules—for example, Safe-RL on humanoids [2020_Javier_safereinforcementlearning], hybrid DRL with identified low-dimensional safety models [2022_Li_UCB_safteystability_model-based_DRL_PPO], footstep-constrained DRL policies [2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation], and reactive DRL steppers operating under feasibility constraints on uneven terrain [2021_DeepQ_DNQ_stepping_animationlonly_steppingstones]. This limitation makes it difficult to prevent the robot from exceeding joint limits, applying excessive torques, or causing self-collisions, particularly when reacting to unexpected events. This is a key area where constrained RL could be applied.
@@
In summary, the pursuit of performance in DRL has often sidelined critical safety issues. The prevalence of blind policies that cannot anticipate environmental hazards, combined with the lack of inherent mechanisms to enforce physical constraints, creates significant risk and hinders real-world deployment. While these challenges are considerable, they also define a clear path forward. The following section on Future Directions and Opportunities explores specific research avenues, such as vision-based learning and safe reinforcement learning, aimed at overcoming these safety barriers and enabling the development of truly robust and reliable bipedal robots.
@@ V-A Pathways for bipedal locomotion
In relation to the research question introduced in Section I, progress in DRL-based bipedal locomotion should be assessed not only through conventional metrics such as reward and success rate but also by broader system-level measures. These include generalisation breadth (across skills, terrains, and morphologies), precision in fidelity-critical tasks (e.g., command-tracking error and foot-placement accuracy), safety and constraint compliance (joint, torque, and contact feasibility), and efficiency or deployability (sample efficiency and on-robot inference latency).
These dimensions build directly upon the challenges outlined in Section IV and together define the key pathways for advancing bipedal locomotion. The following subsections elaborate on these pathways, each addressing one or more of the above aspects to guide progress towards more generalisable and robust control systems.
@@ V-A1 Multi-skill learning
A fundamental goal for the next generation of bipedal robots is to move beyond the paradigm of single-task specialisation and towards versatile skill learning [2023_radosavovic_realworld_transformer_NNmodel, 2024_radosavovic_humanoid_locomotiontokenprediction_transformer, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining]. This research direction focuses on enabling robots to acquire, adapt, and deploy a broad and varied repertoire of motor skills, allowing them to handle unforeseen situations and operate effectively in unstructured environments.
To achieve such versatility, researchers are pursuing several HL pathways, which can be broadly categorised into structured and holistic approaches. The structured approach focuses on explicit decomposition. A prominent example is hierarchical learning, where success depends on appropriately dividing responsibilities; for instance, an HL planner generates reference trajectories, while an LL DRL controller executes them robustly [2024_stanford_humanplus_hierarchical_humanoidshadowingimitation, 2023_template_taskspace_hierarchyscheme_reducedorderstateALIP_learnedhigherlevel_lowlevelinversedynamiccontroller], as shown in Section III. Similarly, skill composition employs a supervisor policy to select and sequence LL experts to solve complex tasks [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task]. A related technique, knowledge distillation, leverages experts by first training them and then distilling their capabilities into a single, compact generalist policy [huang2024diffuseloco].
@@ V-A2 Traversing challenging environments
The goal of versatile skill learning is to enable bipedal robots to traverse challenging, human-centric environments where their unique form offers an advantage. Validating capabilities on such terrains serves a crucial dual purpose. It tests a policy’s generalisation across diverse settings, including stairs and uneven ground, which is essential for real-world integration [2021_siekmann_blind_DRL_stair, 2023_vanmarum_visionDRL_studentteacher_irregularterrain_PPO_periodicrewardfunction, 2023_radosavovic_realworld_transformer_NNmodel]. More critically, it benchmarks precision on treacherous paths such as stepping stones, which demand exact foot placement [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2023_junheng_steppingstones_MPCWBC, 2021_DeepQ_DNQ_stepping_animationlonly_steppingstones]. These environments are the ultimate test of both a robot’s skill repertoire and its control fidelity.
@@ V-A3 Efficient learning and reward shaping
As detailed in Section IV, while DRL has unlocked impressive capabilities in bipedal locomotion, its reliance on training from scratch leads to significant sample inefficiency [2017_DRL_survey, 2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie]. Addressing this bottleneck is a crucial research frontier that calls for both more efficient algorithms and more robust reward designs.
To mitigate sample inefficiency for complex skills, several research pathways are being actively explored. A primary strategy is to leverage prior data rather than learning entirely from scratch. Leveraging prior knowledge provides strong guidance and reduces unsafe exploration by anchoring policies to feasible motion patterns [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2018_xie_OSUDRL_feedback_DRL_reference-based, 2024_stanford_humanplus_hierarchical_humanoidshadowingimitation]. Curriculum learning further organises training from simple to progressively harder tasks, for example standing and balancing before walking and running, which improves stability and convergence [2020_zhaoming_drl_steppingstones_PPOwithactorcritic_referencefree_simulation, 2021_Diego_deepwalk_curriculum_learning_sim2real, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning].
Complementing advances in algorithms is the design of effective and robust rewards. Manual reward engineering remains a significant obstacle, since small choices can induce reward hacking and lengthy tuning cycles [2017_heess_emergence_animation_PPO_comprarisonrlalgorithm_differenttarrains, 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task]. Phase-aware objectives are well established for cyclic gaits such as walking [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie], whereas reward design for non-periodic skills such as jumping is less standardised and often task specific [2023_UCB_Cassie_RL_Jumping]. Promising directions reduce manual effort by adding higher-level guidance, including event-based terms, goal-conditioned objectives, and kinematic reference tracking [2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation, 2021_michael_velocity-basedcontrol_motioncapturedata, 2018_yang_push_recover_wholebody]. Alternatively, learning rewards from data through inverse methods and related approaches aims to replace hand-crafted objectives with implicit ones inferred from demonstrations [2016_GAIL_imitation_learning]. Together, these directions seek to minimise skill-specific tuning and improve the transferability and reliability of learned locomotion policies.
@@ V-A4 Motion retargeting
As human-like agents, bipedal robots—especially humanoids—have the unique advantage of a morphology that is similar to our own. This presents a significant opportunity: the potential to learn from vast libraries of human motion data. While large-scale datasets such as AMASS [2019_Naureen_AMASS] and Motion-X [2024_motion_X_datasets] provide a wealth of such data, they are inherently human-centric and cannot be used directly, requiring substantial retargeting effort [2024_speratebody_expressive_allmotion]. Therefore, motion retargeting emerges as a critical component to bridge this gap. The challenge of this pathway is not merely to transfer human movements to the robot, but to generate trajectories that are both high in stylistic fidelity and physically feasible, adhering to the robot’s unique dynamics and constraints. Successfully developing these retargeting methods provides a scalable solution for accessing the data needed to train the natural and versatile generalist policies of the future.
@@ V-A5 Bridging the gap from simulation to reality
Strategies to bridge the sim-to-real gap generally follow two main philosophies. The first aims to train policies robust enough to tolerate the inevitable mismatch between simulation and reality, while the second focuses on minimising the gap itself by making the simulator a more faithful replica of the physical world.
The first approach seeks to reduce the discrepancy by improving the simulation’s fidelity. This is often achieved through system identification (SI), where real-world robot data are used to fine-tune the simulator’s parameters to create a more accurate “digital twin” [2019_Yu_Phy_system_identification, 2023_masuda_simtoreal_identification]. This can include explicitly learning complex actuator dynamics to model the motors’ behaviour [2019_Yu_Phy_system_identification, 2019_simtoreal_actuatornet_quadrupedrobot]. Other methods, such as designing specialised feedback controllers [2021_Casillo_cyberbotics_hierarchyscheme_bezierpolynominals], also contribute by making the system less sensitive to residual modelling errors.
In contrast, the second philosophy accepts that simulations will always be imperfect and instead focuses on creating highly adaptive, robust policies. The primary method here is DR, which forces a policy to generalize by training it across a wide range of simulated physical variations. Other various ways, such as through end-to-end training that uses measurement histories to adapt online like in RMA [2022_Ashish_RMA_sim2real_UCB], or via policy distillation, where a privileged “teacher” guides a “student” policy [2023_vanmarum_visionDRL_studentteacher_irregularterrain_PPO_periodicrewardfunction] to have a knowledge of unknown information like friction . Additionally, techniques like adversarial motion priors [2024_Zhang_wholebody_adversarial_motion_priors, 2024_tang_humanmimic_adversarial] are used to ensure the learned behaviours are not just robust but also physically plausible.
@@ V-A6 Perception-conditioned locomotion
Integrating exteroceptive sensors such as cameras and LiDAR enables bipedal robots to proactively plan footsteps, avoid obstacles, and adapt to upcoming terrain. This shift from reactive to anticipatory control is essential for navigating unstructured real-world environments.
The vision-based pathway is a human-inspired approach using RGB and depth cameras to capture rich data on colour, texture, and object appearance [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2023_duan_OSUDRL_heightmap_visionbased_hybrid, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning]. In contrast, LiDAR is an active sensing method that generates precise 3D point clouds of the terrain. While vision provides richer data but is sensitive to lighting, LiDAR offers robust geometric measurements without visual detail.
Based on this sensory data, current research is exploring two primary pathways for processing perceptual information for control. The first involves creating an intermediate geometric representation, such as a height map from scanners [2023_duan_OSUDRL_heightmap_visionbased_hybrid]. This provides the policy with structured topographical data for effective footstep planning. The second is a more end-to-end approach, which utilises direct vision inputs such as RGB or depth images as inputs to the RL policy for real-time decision-making [2018_Tsunekawa_visualnavigation_hierarchy_DDPG-sim2real_NAO, 2022_byravan_nerf2real_basketball]. The former offers interpretability, while the latter promises more nuanced, reactive behaviours learned directly from raw perception.